x 2 +2x+1=0,x=?

x2+2x+1=0,x=?
 

Grade:10

4 Answers

RAJU
8 years ago
x=-1
OUTPASS
8 years ago
-1=x.
AMMANANNA
8 years ago
for this question is the expansion is
x^2+2x+1=0
for u have put x= -1 in that equqtion  you will get 0 so x= -1 value is satisfies the equation
so that the answer is x= -1
thank u boss
RAMCHANDRARAO
8 years ago
roots for the equation ( x2+2x+1) is x=-1
this equation also written as (x+1)^2=0 that implies x+1=0
therefore x=-1